Ashley has $120, 000 to invest and decides to put some in a CD that earns 4% interest per year and the rest in a low-risk stock that earns 3%. How much did she invest in each to earn $ 3800 interest in the first year?
Found 2 solutions by solver91311, greenestamps:
Answer by solver91311(24713)





Solve for and then calculate

John

My calculator said it, I believe it, that settles it


Answer by greenestamps(13195)



If a formal algebraic solution is not required, here is a non-algebraic solution method that will get you to the answer much faster and with less effort than algebra.

All $120,000 invested at 3% would yield $3600 interest; all invested at 4% would yield $4800 interest.

The actual interest, $3800, is $200/$1200 = 1/6 of the way from $3600 to $4800.

That means 1/6 of the total was invested at the higher rate.

ANSWER: 1/6 of $120,000, or $20,000 was invested at 4%; the other $100,000 at 3%.

CHECK: .04(20,000)+.03(100,000) = 800+3000 = 3800